# AtomSurf: Surface Representation for Learning on Protein Structures

### v3

@@ Appendix E Validation dataset description
This task aims to segment 5s ribosomal RNA molecules into functional components, as defined by aligned regions in a Multi-Sequence Alignment Poulenard et al. (2019). This dataset consists of 640 RNA surface meshes of about 15k vertices. It was introduced as a benchmark to compare surface encoders.
@@
This task aims to predict which part of a protein interacts with which part of another. Framed as a classification task, pairs of residues from two proteins are labeled as positives if they interact and as negatives if they do not. The dataset comprises 87k, 31k, and 15k training, validation, and test examples.
@@
The objective here is to determine if a mutation enhances the stability of protein-protein interaction. Given a protein-protein interaction structure and its mutated version, this classification task labels the pair as a positive example if it exhibits increased stability. This task includes 2864, 937, and 347 examples in each data split. For both PIP and MSP, the performance metric is the Area under the Receiver Operating Characteristic curve (AuROC).
@@
PSR is a regression task and aims to assign a quality score to predicted protein structures from the Critical Assessment of Methods of Protein Structure Prediction (CASP) Kryshtafovych et al. (2019) competition. The PSR data train, validation, and test splits hold 25.4k, 2.8k, and 16k systems respectively. The "global $R_{S}$" term represents the mean correlation across all systems and proposals. Meanwhile, "mean $R_{S}$" refers to the average correlation for each system.
Masif-ligand: In addition to the Atom3d benchmark tasks, we evaluate our proposed approach on the task of ligand-binding preference prediction for protein binding sites, introduced in Gainza et al. (2020). Binding sites are regions of protein surfaces where small molecules (ligands) interact with the protein. A fine characterization of those regions helps understanding protein functions and plays a significant role in drug design and discovery. This dataset consists of protein binding sites that accommodate one of seven co-factors. It comprises 1,634 training instances, 202 validation instances, and 418 test cases. Given a binding site, the task amounts to predicting its corresponding co-factor.
@@ Appendix F Supplementary results
F.1 Performance of the proposed Surface Encoder
We validate our surface encoder on the task of segmenting 5s ribosomal RNA molecules into functional components, as defined by aligned regions in a Multi-Sequence Alignment Poulenard et al. (2019). This dataset consists of 640 RNA surface meshes of about 15k vertices, and was already used to compare modern surface encoders.
We assess the impact of the proposed enhancements to DiffusionNet by showing the learning curves of the enhanced models on the RNA segmentation task (see Figure 13). Moreover, we compare their performance to standard recent surface encoders, DGCNN Wang et al. (2019) and DeltaConvWiersma et al. (2022). Our experiments show that our adjustments significantly improve the network’s stability issues and allow it to converge to a model with better performance than other networks.
@@ F.1 Performance of the proposed Surface Encoder
Figure 12: Evolution of the learning curve (loss) on the RNA segmentation task using the original and our enhanced DiffusionNet models. It is evident that the original implementation is unstable and exhibits peaks in the loss.
Figure 13: Performance of Different Surface Encoders on the RNA Segmentation Task.
F.2 Validation of the new model on PSR
We conduct a similar analysis of the proposed enhancements on the PSR task. We depict the obtained learning curves in Figure 15 and report performance in Figure 15. Again, we observe that our suggested enhancements result in a more stable learning.
@@ F.2 Validation of the new model on PSR
Figure 14: Global correlation learning curve on the PSR task before and after applying our enhancements to DiffusionNet.
[…]
[0.54]
PSR
Local R Global R
DiffusionNet (original) 0.17 0.39
DiffusionNet (enhanced) 0.33 0.643
Figure 15: Results of our enhanced model on the PSR task.
F.3 Performance on binding site detection
Below are the results for the binding site detection task, in the format of a table for exact comparisons.
Table 4: Performance of our hybrid approach on AbAG task.
Antibodies Antigens
AuROC MCC AuROC MCC
GEP (Pegoraro et al., 2024) 0.80 0.28 0.72 0.18
AtomSurf 0.86 0.47 0.73 0.18
F.4 Assessment of the impact of the threshold distance
@@
Table 5: Performance on the PSR task as a function of the threshold used.

### v4

@@ Appendix E Validation dataset description
This task aims to segment 5s ribosomal RNA molecules into functional components, as defined by aligned regions in a Multi-Sequence Alignment (Poulenard et al., 2019). This dataset consists of 640 RNA surface meshes of about 15k vertices, split at random. It was introduced as a benchmark to compare surface encoders.
@@
This task aims to predict which part of a protein interacts with which part of another. Framed as a classification task, pairs of residues from two proteins are labeled as positives if they interact and as negatives if they do not. The dataset comprises 87k, 31k, and 15k training, validation, and test examples, split based on a 30% sequence identity. Moreover, proteins in the test set are presented in their apo conformation.
@@
The objective here is to determine if a mutation enhances the stability of protein-protein interaction. Given a protein-protein interaction structure and its mutated version, this classification task labels the pair as a positive example if it exhibits increased stability. This task includes 2864, 937, and 347 examples in each data split, and the splitting is performed based on a 30% sequence identity. For both PIP and MSP, the performance metric is the Area under the Receiver Operating Characteristic curve (AuROC).
@@
PSR is a regression task and aims to assign a quality score to predicted protein structures from the Critical Assessment of Methods of Protein Structure Prediction (CASP) (Kryshtafovych et al., 2019) competition. The PSR data train, validation, and test splits hold 25.4k, 2.8k, and 16k systems respectively. Splits correspond to a time split, with more recent CASP competition belonging to the test split. The $R_{g}$" term represents the mean correlation across all systems and proposals. Meanwhile, $R_{l}$" refers to the average correlation for each system.
Masif-ligand: This task aims to predict ligand-binding preferences for protein binding sites, as introduced in (Gainza et al., 2020). Binding sites are regions of protein surfaces where small molecules (ligands) interact with the protein. A fine characterization of those regions helps understanding protein functions and plays a significant role in drug design and discovery. This dataset consists of protein binding sites that accommodate one of seven co-factors. It comprises 1,634 training instances, 202 validation instances, and 418 test cases, split using sequence identity. Given a binding site, the task amounts to predicting its corresponding co-factor.
AbAg: This dataset focuses on binding site prediction in the context of immunology: predicting the interaction between an antibody and its target, denoted as the antigen. Improved understanding and ability to predict such interactions has direct applications in antibody-based treatments. (Pegoraro et al., 2024) introduces a dataset containing 235 antibody–antigen complexes, with 186 for training and 49 for testing, split based on a 70% similarity threshold. The task consists in predicting the residues involved in the binding site on the antibody and antigen separately. The performance on the antibody side is computed only on the Complementarity Determining Regions (CDRs). They propose different methods and we chose to compare against MIXmean-egnn PiNet that had the most balanced performance.
PINDER: This is a recent, large-scale dataset (Kovtun et al., 2024) that holds over 2M interaction in its full version. We used the clustered version of this dataset that holds around 42k structures and reduces redundancy. Pinder also introduces a strict splitting criterion based on interface similarity, that was shown to reduce leakage. In addition to their bound form (holo), systems in the test set are also available in their unbound form (apo) and as predicted by AlphaFold (predicted), representing a more realistic use case. The number of systems present in the test set is 1955 clusters, with 342 corresponding apo and 1747 predicted structures. This is two orders of magnitude higher than AbAg.
We introduce two tasks related to binding site prediction. Our first task is formulated as PIP (Pinder-Pair): given two interacting monomers taken in isolation, it aims at classifying pairs of residues into interacting, and non-interacting. The other task is close to Masif-Site (Gainza et al., 2020), only taking one protein as input to predict interacting residues (Pinder-site).
@@ Appendix F Supplementary results
F.1 Validation of the new model on PSR
We conduct a similar analysis of the proposed enhancements to the one on ribosomal RNA on the PSR task. We depict the obtained learning curves in Figure˜11 and report performance in Figure˜11. Again, we observe that our suggested enhancements result in a more stable learning.
@@ F.1 Validation of the new model on PSR
F.2 Variation of our results over different seeds
In table 3, we computed the standard error of the mean computed over three replicates. The standard error is small compared to the gap of performance with the next best method, hinting for a statistically significant difference (unfortunately, there are no replicates for other methods to perform actual statistical tests). As expected, the standard error is higher for smaller datasets (MSP, AbAg) than larger ones (PIP, PSR).
Supplementary Table 3: Standard error of our model evaluated on three replicates for each of the task included. As a reference, we include the best performing competing model, without pretraining.
Task PIP MSP PSR
Metric AuROC AuROC $R_{l}$ $R_{g}$
Runner-up method 87.1 68.5 63.2 86.8
Ours - Mean 90.9 71.5 61.7 85.7
Ours - standard error 0.088 1 0.27 0.22
Task Ab Ag Masif Ligand
Metric AuROC MCC AuROC MCC Balanced acc.
Runner-up method 80 28 72 18 81
Ours - Mean 88 53 67.3 14.3 88.2
Ours - standard error 1 2.1 1.8 1.5 0.3
F.3 Performance on PINDER task
Table˜4 shows the performance of AtomSurf, compared to ProNet on the clustered PINDER dataset. In this dataset, the systems are clustered to remove redundancy. Moreover, the splitting strategy is stricter, based on a similarity score of the interfaces. In addition to the canonical test set, the test set systems are also available in their unbound form (apo setting) and as predicted by AlphaFold (predicted setting). This results in a more realistic estimate of the quality of the predictions. We report the accuracy and AuROC on both tasks, on each of those splits. Finally, we also report its performance without ESM embeddings.
Task Pinder-Pair
Split Holo Apo AF2
Metric Auroc Acc Auroc Acc Auroc Acc
Pronet 80.1 72.5 78.2 71.5 73.5 67.9
AtomSurf 92.8 85.3 88.4 80.7 87.1 80.2
AtomSurf - no ESM 93 85.4 87.6 80.2 87 79.9
Task Pinder-Site
Split Holo Apo AF2
Metric Auroc Acc Auroc Acc Auroc Acc
Pronet 74.3 69.2 70.7 71 60.6 56.3
AtomSurf 88.3 82.3 84.2 80.8 82 76.6
AtomSurf - no ESM 87.6 79.2 82.4 77.1 80.9 78.8
Supplementary Table 4: Performance on the Pinder Task.
F.4 Model ablation
Here, we present the detailed ablation study of our method AtomSurf-bench. We examine the impact of different design choices on our tasks. All of these results are obtained in benchmark settings, both in terms of features (only atom-type as input) and of surface encoders and graph encoders (GCNs).
First, we look at the overall architecture of our approach, following Appendix˜C, where we introduce the sequential and bipartite scenarios, which amount to variations in the connectivity of the blocks. In particular, the HoloProt method proposed in Somnath et al. (2021) falls into the sequential framework with just one encoding block and a MeshCNN (Hanocka et al., 2019) for protein encoder. For fairness, we replace MeshCNN with our encoder and refer to this setup as HoloProt.
Another major design choice is the choice of the Message Passing (MP) component. We explore the use of three possible message-passing networks. Motivated by the success of DGCNN (Wang et al., 2019), in our Att. setting, we discard the geometric notion of a neighborhood, allowing for potentially long-distance message passing. In this setting, all nodes from the graphs attend to all vertices from the surface. To deal with the incurred computational burden, we use the recent memory-efficient Flash Attention (Dao et al., 2022). We also explore the use of more conventional Graph Convolutional Networks (GCN setting) (Kipf & Welling, 2016) and the use of Graph ATtention networks (Veličković et al., 2017; Brody et al., 2021) for our final GAT setting. Finally, we also try using three or four blocks in our networks, always adjusting the network width to keep the number of parameters constant. Our results are presented in Table˜5.
Supplementary Table 5: Ablation study of our method: We compare various architectural designs, and message-passing methods for our task.
MSP PIP PSR
Method MP AuROC AuROC local R global R
sequential - 0.609 0.855 0.319 0.71
HoloProt - 0.537 0.824 0.383 0.715
Att. 0.689 0.791 0.402 0.792
bipartite GCN 0.697 0.868 0.421 0.797
GAT 0.707 0.876 0.452 0.833
F.5 ESM embeddings ablation
We conducted an Ablation of our method by training models without the ESM embeddings as input features. We do not change hyperparameters of our method to make up for this change, which strongly disrupts the training on smaller datasets, and especially on MSP which fails to train. We report the result in Table˜6 on all tasks but Pinder, and in Table˜4 for the PINDER task.
Supplementary Table 6: Performance of our model when ablating the ESM node embeddings.
Task PIP MSP PSR
Metric AuROC AuROC $R_{l}$ $R_{g}$
Runner-up method 87.1 68.5 63.2 86.8
Ours - Mean 90.9 71.5 61.7 85.7
Ours - no ESM 90 54.5 47 82
Task Ab Ag Masif Ligand
Metric AuROC MCC AuROC MCC Balanced acc.
Runner-up method 80 28 72 18 81
Ours - Mean 88 53 67.3 14.3 88.2
Ours - no ESM 86 48 68 14.5 84
The performance decreases as expected, showing that ESM is truly an important feature. However, even without ESM, our method still shows significant results on most tasks, such as PIP, AbAg and MasifLigand. The overall difference was more noticeable on smaller datasets, especially on MSP for which the ablated network failed to learn a useful signal without ESM. Considering that AtomSurf-bench (which does not use ESM embeddings) gets 70.9 AuROC on MSP, we believe a high performance could be recovered by tuning the optimization. Based on this ablation, we conclude that ESM features are an important source of information, but that even without them, integrating a surface and a graph method gives a strong performance.
On PINDER, we obtain a small benefit from using ESM embeddings as node features, most notably on the apo set. However, this gap seems to be quite small on this dataset with stricter splits. One possible explanation for this result would be that those sequence-derived features contribute to memorization, whilst learnt structural properties could better generalize.
F.6 Assessment of the impact of the threshold distance
@@
Supplementary Table 7: Performance on the PSR task as a function of the threshold used.